OpenStudy (anonymous):

How would I solve this (help would be greatly appreciated :) ) (sqrt(-2x))-(sqrt5-x)=-3/(sqrt5-x)

OpenStudy (anonymous):

\[\sqrt(-2x)-(\sqrt5-x)=-3/(\sqrt(5-x))\]

OpenStudy (anonymous):

Is this what you have? \[\sqrt{-2x}-\left( \sqrt{5}-x \right)=\frac{ 3 }{ \sqrt{5}-x }\]

OpenStudy (anonymous):

the square root goes over (5-x) not just 5

OpenStudy (anonymous):

In both?

OpenStudy (anonymous):

\[\sqrt{-2x}-\sqrt{5-x}=\frac{ 3 }{ \sqrt{5-x} }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, what if we multiply both sides by \[\sqrt{5-x}\]

OpenStudy (anonymous):

on the right, we will be left with (-3) (Note, I missed the negative)

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

On the left we have to distribute so we will mult \[\sqrt{5-x}\] by \[\sqrt{-2x}\] and by \[\sqrt{5-x}\]

OpenStudy (anonymous):

we get:\[\sqrt{-2x \left( 5-x \right)}-\left( 5-x \right)=-3\]

OpenStudy (anonymous):

please do look for errors on my part now, clean up - clear those ( )'s by distributing the neg and mult the expressions in the radical once all is cleaned up, GET THE RADICAL ALONE ON ONE SIDE

OpenStudy (anonymous):

how are we doing?

OpenStudy (anonymous):

Good! just doing the distribution. I've got a question though... where did the square root over the -5-x go?

OpenStudy (anonymous):

\[\sqrt{5-x} \sqrt{5-x}=5-x\] It is like squaring a square root - they are opposite operations and undo each other

OpenStudy (anonymous):

so we get -(5 - x) which is - 5 + x

OpenStudy (anonymous):

oh, okay!

OpenStudy (anonymous):

Cpuld you walk me through further?

OpenStudy (anonymous):

I've got